OpenStudy (anonymous):

What are 3 solutions for the equation. 2y=3x+11

OpenStudy (precal):

this equation has many solutions, they are asking you to name at least 3 just pick randon points try x=0 y=? x=1 y=? x=2 y=? just sub those x values one by one and calculate the y values, I can check them for you if you like

OpenStudy (anonymous):

x=3 y=1

OpenStudy (anonymous):

@precal

OpenStudy (anonymous):

well is it y=3 and x =10?

OpenStudy (precal):

if x = 3 then 2y=3x+11 2y=3(3)+11 2y=9+11 2y=20 y=10

OpenStudy (precal):

sorry I made a small mistake earlier, now I fixed it

OpenStudy (anonymous):

Ok i thought it was correct haha.

OpenStudy (precal):

ok so you have one point (3,10) go find 2 others

OpenStudy (precal):

let me know what they are and I will check them for you